OpenStudy (anonymous):

Help I will fan and Metal! Which is an equation for the line with slope 1/3 through (-2,1)? A.)y=1/3x+1 B.)y=1/3x+5/3 C.)y=1/3x-5/3 D.)y=1/3x+1/3

OpenStudy (anonymous):

@SolomonZelman ? @satellite73 ???

OpenStudy (anonymous):

Anyone?

OpenStudy (solomonzelman):

you know the slope is 1/3, so \(\large\color{black}{ y=\frac{1}{3}x+b }\)

OpenStudy (anonymous):

Is it A? y=1/3x+1? @SolomonZelman

OpenStudy (solomonzelman):

B

OpenStudy (anonymous):

Oh okay.